LOCAL LINEARIZATION METHOD FOR THE NUMERICAL SOLUTION OF STOCHASTIC DIFFERENTIAL EQUATIONS

R. BISCAY, J. C. JIMENEZ, J. J. RIERA AND P. A. VALDES

Centro de Neurociencias de Cuba, Apartado 6880, Ciudad de la Habana, Cuba

(Received October 17, 1994; revised September 27, 1995)

Abstract. The Local Linearization (LL) approach for the numerical solution of stochastic differential equations (SDEs) is extended to general scalar SDEs, as well as to non-autonomous multidimensional SDEs with additive noise. In case of autonomous SDEs, the derivation of the method introduced gives theoretical support to one of the previously proposed variants of the LL approach. Some numerical examples are given to demonstrate the practical performance of the method.

Key words and phrases: Stochastic differential equations, numerical solution, local linearization.